# FLTrojan: Privacy Leakage Attacks against Federated Language Models through Selective Weight Tampering

### v3

@@ 5.1.1 Dataset Description
Our experiments include both canaries and real-world sensitive data. To observe how dataset size affects attack performance, we selected two datasets: the Penn Treebank (PTB) dataset [43], which is a small dataset consisting of only 5MB of text from financial news articles, and the Wikitext-103 (Wiki) dataset [47], which is a comparatively larger dataset consisting of 181MB of text. Similar to [8], we augment both of these datasets with out-of-distribution artificial canaries, which we manually crafted instead of web scraping to ensure that the pre-trained LLMs have not seen these texts before in their pre-training phase. These canaries contain dummy privacy-sensitive information. We insert four types of private information as canaries: phone number, credit card number, email, and physical address. Some examples of the canaries are provided in the Appendix (Table 6). It is worth mentioning that the training dataset may contain out-of-distribution texts that are not privacy-sensitive, such as product keys and hotline numbers. However, in our attacks, the adversary specifically designs the prompts to extract only the privacy-sensitive information they are interested in. The presence of other non-privacy-sensitive outlier texts in the dataset does not affect the effectiveness of the attack.
@@
To confirm that our proposed method does not merely work for artificially inserted canaries, we further include a dataset in our experiments that already has several naturally occurring sensitive information. The Enron email dataset [36] is a well-known collection of emails that were exchanged among employees of the Enron Corporation, an American energy company. It contains private information, e.g., phone numbers, addresses, login credentials, and potentially other sensitive details about individuals involved in the company. We only retain the body of the emails and pre-process them by filtering out all attachments and auto-generated emails. We take a subset of the resultant dataset, which is 412MB in size, and we manually identify 125 privacy-sensitive sequences from it. Several existing works [8, 42, 24] on LLM’s privacy conducted their experiments with these three datasets.
@@
For the Enron email dataset, we observe that the majority of the 125 private sequences have multiple occurrences. Approximately 29% of them occur twice, 18% occur three times, and 7% occur four or more times. This implies that although state-of-the-art defenses like differential privacy [16] provide guarantees under the assumption that records are unlikely to be duplicated, this may not be true for real-life datasets. To simulate this phenomenon in the PTB and Wikitext datasets, we generate 200 unique canaries and insert them into the datasets with $k$ repetitions. Specifically, we consider $k\in{\{1,10,50\}}$, meaning one canary per $\sim{\{46K,4.6K,920,460\}}$ training sequences for PTB, and one canary per $\sim{\{3.5M,350K,70K,35K\}}$ training sequences for Wikitext.
5.1.2 Training and Attack Implementation
We describe the FL setup, attack settings, and evaluation criteria as follows.
@@ 5.1.2 Training and Attack Implementation
FL Setup: We fine-tune Gemma-2b [65], Llama-2 [68], GPT-2 base [59] and BERT [13] as representative language models (LM) to evaluate our proposed attacks. Later we also observe the changes in attack performances for the medium and large versions of GPT-2. Due to computation resource limitations, we could not train the Llama-2 model in a large-scale FL setting. Instead, we conducted our attacks on Llama-2 in a compact experimental setup. The relevant experiments with Llama-2 are included in Appendix E. Additional information about these models is available in the Appendix B. For the primary FL task, we have 100 participants and randomly select 20 for each round. We vary the number of attackers and victims between 1 and 10 and distribute the sensitive sequences among the victims accordingly. We train  [^2] Experiments are run on an 8-core NVIDIA RTX A6000 GPU Server the model for 700 iterations, each with five local epochs, using an initial learning rate of $10^{-4}$ with a linearly decayed learning rate scheduler.
@@
❏ We use a dataset of English plaintext jokes [56] as $\mathcal{D}_{reg}$ as discussed in Sections 3 and 4. This dataset exclusively contains $1622$ regular English sentences without privacy-sensitive sentences. Besides, we generate $100$ dummy privacy-sensitive sentences to use as $\mathcal{D}_{sen}.$ As mentioned earlier, it does not overlap with the original private texts. The attacker uses their prior knowledge of the sensitive data pattern to generate these dummy sentences and capture the model’s behavior towards out-of-distribution data. The number of epochs for finetuning $G_{r}$ and $G_{s}$ was empirically set to 20 with early stopping enabled.
@@
❏ We use the Adam [34] optimizer in the selective weights optimization (SWO) algorithm. We empirically tune the number of steps between 200 and 1000, the clipping threshold between 1 and 4, and set the initial learning rate at $1\times 10^{-3}$.
❏ Since model weights do not fall into conventional categories of data, we experiment with several learning algorithms for the transformation task to identify the most suitable one. These algorithms include normalizing flows (e.g., planer flows [38] and RealNVP[14]), dense autoencoder (DAE), variational autoencoder (VAE)[35], convolutional autoencoder (CAE)[79], and transformer[69]. We find that DAE, CAE, and transformer showed comparable performances (in terms of validation loss) while normalizing flows and VAE did not yield promising outcomes. However, we eventually chose to employ DAE for this task as the other two, particularly the transformer, turned out to be computationally much more expensive taking roughly twice the time of DAE to complete the training.
❏ To ensure better stability during the transformation, we use Z-score normalization [2] to normalize the weight vector before passing it to the SWO or WTL algorithm. After the transformation, we denormalize the output weight to its original scale. For the sake of simplicity, we skip the normalization step in Algorithms 2 and 3.
Evaluation Criteria: For the autoregressive text generation, we consider every sensitive sequence $x=p||c$ in the model’s training dataset, where x is split into a prefix $p$ (i.e., the non-sensitive component) and a suffix $c$ (i.e., the sensitive component). The suffix is the attacker’s point of interest. For each sequence, we report if the model exactly reproduces $c$ when prompted with $p$, following the data reconstruction procedure described in Section 4.5.1. In masked language modeling, for every sensitive sequence $x=p||c||s$ in the training dataset, we aim to reconstruct the masked-out sensitive component $c$, given a prefix $p$ and suffix $s$. We also evaluate the sentence-level membership inference [48] where the attacker attempts to infer whether a sensitive sequence $x$ is part of the victim’s local training data or not.
5.2 Baselines and Metrics
@@
Intermediate Rounds (IR): The attacker has access to the ground truth information of intermediate victim rounds from server cooperation as mentioned in Section 4.3. IR directly prompts these victim models without applying any attack strategy.
@@
Metrics. We report the number of successful reconstructions (NSR) of the artificial canaries (for Wikitext and PTB) and the inherent sensitive texts (for Enron). We also evaluate the level of leakage using exposure (EXP), a metric introduced in [8], where higher exposure indicates greater privacy leakage. For membership inference, we report the recall and precision following [48]. Moreover, we use validation perplexity (PPX) as a metric for the performance of the model, where lower perplexity refers to better model utility.
Figure 5: Number of successful reconstructions (NSR) out of 200 unique canaries for different attack strategies along with the baselines for (a) Wikitext with Gemma, (b) Wikitext with GPT-2, (c) Wikitext with BERT. Then, out of 125 in-house sensitive sequences for (d) Enron with Gemma, (e) Enron with GPT-2, and (f) Enron with BERT. Finally, membership inference (g) recall and (h) precision scores for different attack strategies along with three baselines on the Wikitext Dataset using BERT.
5.3 Research Questions
We aim to answer the following research questions through a comprehensive empirical study.
@@ 5.3 Research Questions
RQ 4: How does the inclusion of the server as an adversary impact the attack performance?
5.4 Results Summary
@@
Table 2: Recall (%) and precision (%) of victim round identification step from 700 rounds.
5.4.1 Victim Round Identification Performance
Table 2 demonstrates the results of our victim round identification mechanism based on 700 FL rounds for both recall and precision. It yields high recall scores in all cases, suggesting that it can identify the victim rounds with significant success, which answers RQ 1. Besides, we can see a growth in recall in the dynamic attack and after increasing number of victims. This is because, in both cases, victim model updates are considered more frequently during training, bolstering the global model’s memorization of the victim’s data, which enhances the VRI process. However, those extra victim updates during training impact non-victim training rounds as well, which generates more false-positive outputs. Hence, with the increased number of victims, the precision scores fall slightly (confusion matrices in Figure 12).
5.4.2 Data Reconstruction Performance
Figure 5(b) and 5(a) demonstrate the results across different attack strategies and the four baselines on the Wikitext dataset, in terms of the number of successful reconstructions (NSR) when autoregressive text generation is performed with Gemma and GPT-2 respectively. Overall, The Gemma model achieves a superior NSR compared to GPT-2, mainly due to its larger size and complexity. First, comparing the scores of the first two baselines answers RQ 2. A higher score of IR compared to LR indicates that model snapshots from intermediate victim rounds cause greater leakage than only the final model (hypothesis 2), which justifies the importance of our victim round identification step. Next, for both Gemma and GPT-2, our static attack outperforms IR by 15-18% with both SWO and WTL, answering RQ 3. This shows that it is indeed possible to significantly increase the LM’s memorization of clients’ private data by maliciously tampering with its weights. Additionally, WTL performs better than SWO in all cases, proving it to be a better MDM technique. Apart from that, introducing active threats during training further increases privacy leakage, as our dynamic mode of the attack has a much better NSR (137 for Gemma) than the static mode (125). Furthermore, incorporating victims’ local updates in place of their global counterparts with cooperation from the server boosts the attack performance considerably (by 6 to 8%) in both $static^{+}$ and $dynamic^{+}$ modes of the attack. In fact, $dynamic^{+}$ is our best-performing attack variant with $71\%$ (NSR=142) successful reconstructions.
@@ 5.4.2 Data Reconstruction Performance
Figure 5(c) illustrates attack performance in masked language modeling, i.e., with BERT [13]. Overall, NSR scores for BERT are much lower than GPT-2. The improvement in attack performance for WTL over SWO is also not as prominent in this case as it was for GPT-2. Jagannatha et al. [28] show that masked LMs have lower privacy leakage than autoregressive LMs. This could be attributed to the difference in the architecture of these two types of LMs and how they understand the context. BERT understands the context from both the left and the right sides of a masked-out word, whereas Gemma/ GPT-2 uses only the left context to predict the next word. BERT’s bidirectional attention might help it distribute its focus, thereby reducing the chances of overfitting to specific examples. However, our best attack strategy without any server cooperation (dynamic) derives around $48\%$ (95 out of 200) successful reconstructions, while with $dynamic^{+}$the accuracy reaches $55\%$. From the results, it is quite evident that our proposed attack methods are highly effective against masked language modeling in addition to autoregressive text generation. Due to space constraints, we put the experiment results with the PTB dataset in Appendix F.
@@
Finally, the performance of our attacks on real-world data, i.e., the Enron email dataset, is depicted in Figures 5(e) and 5(f). It is worth mentioning that the privacy-sensitive sequences in this dataset are less structured than the artificial canaries, making the attacks more challenging and having some interesting impact on the results. Firstly, both the static and $static^{+}$ attacks slightly underperform in this scenario compared to artificial canary extraction. As shown in Figures 5(e) and 5(f), the NSR scores of the static attack for both Gemma/GPT-2 and BERT are quite close to IR’s score. Nevertheless, active interference by the adversary changes the picture notably, as we can see the dynamic attack on Gemma increases the NSR by around 10% from the static variant, while the $dynamic^{+}$ attack manages to extract 64% (80 out of 125) of the private sequences. The results show that our attacks can successfully compromise the privacy of real-world data with high accuracy. Apart from that, similar to the Wikitext dataset with canaries, masked language modeling with BERT turns out to be more resilient than Gemma/GPT-2 against leakage for the Enron dataset with real-world private data (Figure 5(f)). Results with respect to the exposure (EXP) metric are moved to the Appendix (D).
5.4.3 Membership Inference Performance
@@
We adopt the implementation of [48] for membership inference based on the likelihood ratio hypothesis, where part of our test set consists of all the canaries (for PTB and Wikitext) or private texts (for Enron). We insert 200 canary-like false-positive (outside training data) samples into the test set, which are semantically close to the canaries but do not match any of their secret portions. The evaluation is done over 400 samples (200 canaries + 200 false positives) for both Wikitext and PTB. For the Enron dataset, we add 125 false positives into the test set along with the 125 already identified private sequences, making 250 samples in total. Looking at the results for Wikitext in Figure 5(g), we can see that the static mode significantly improves the recall score (by 15-26%) compared to the lower baselines (LR, IR). This demonstrates the effectiveness of our MDM algorithms in identifying the canaries and confidently inferring their membership. We achieve the highest recall of 61% with the $dynamic^{+}$ version, which shows results similar to the upper baselines (FP and MU). Our MDM approach directs the language model to prioritize sensitive training samples, leading to assigning a higher likelihood to those samples. This reduces the number of missed positive samples, increasing the true positive count and decreasing the false negative count. Consequently, the recall rate experiences a significant improvement due to these dual effects. However, there was only a slight improvement in the ratio of true positives to false positives. As a result, the attack performance did not fluctuate much across different attacks and the baselines in terms of precision as shown in 5(h). For a better understanding of the phenomenon, please refer to the confusion matrices in Figure 12 and Section G.3 of the Appendix. We also put the membership inference results for PTB and Enron into the Appendix (Section G.2) for space constraints.
5.5 Attack Cost Analysis

### v4

@@ V-A Dataset Description
Our experiments include both canaries and real-world sensitive data. To observe how dataset size affects attack performance, we selected two datasets: the Penn Treebank (PTB) dataset [marcus1993building], which is a small dataset consisting of only 5MB of text from financial news articles, and the Wikitext-103 (Wiki) dataset [merity2016pointer], which is a comparatively larger dataset consisting of 181MB of text. Similar to [carlini2019secret], we augment both of these datasets with out-of-distribution artificial canaries, which we manually crafted instead of web scraping to ensure that the pre-trained LLMs have not seen these texts before in their pre-training phase. These canaries contain dummy privacy-sensitive information. We insert four types of private information as canaries: phone number, credit card number, email, and physical address. Some examples of the canaries are provided in the Appendix (Table VI). It is worth mentioning that the training dataset may contain out-of-distribution texts that are not privacy-sensitive, such as product keys and hotline numbers. However, in our attacks, the adversary specifically designs the prompts to extract only the privacy-sensitive information they are interested in. The presence of other non-privacy-sensitive outlier texts in the dataset does not affect the effectiveness of the attack.
@@
To confirm that our proposed method does not merely work for artificially inserted canaries, we further include a dataset in our experiments that already has several naturally occurring sensitive information. The Enron email dataset [klimt2004introducing] is a well-known collection of emails that were exchanged among employees of the Enron Corporation, an American energy company. It contains private information, e.g., phone numbers, addresses, login credentials, and potentially other sensitive details about individuals involved in the company. We only retain the body of the emails and pre-process them by filtering out all attachments and auto-generated emails. We take a subset of the resultant dataset, which is 412MB in size, and we manually identify 125 privacy-sensitive sequences from it. Several existing works [carlini2019secret, lukas2023analyzing, gupta2022recovering] on LLM’s privacy conducted their experiments with these three datasets.
@@
For the Enron email dataset, we observe that the majority of the 125 private sequences have multiple occurrences. Approximately 29% of them occur twice, 18% occur three times, and 7% occur four or more times. This implies that although state-of-the-art defenses like differential privacy [dwork_roth_dp] provide guarantees under the assumption that records are unlikely to be duplicated, this may not be true for real-life datasets. To simulate this phenomenon in the PTB and Wikitext datasets, we generate 200 unique canaries and insert them into the datasets with $k$ repetitions. Specifically, we consider $k\in{\{1,10,50\}}$, meaning one canary per $\sim{\{46K,4.6K,920,460\}}$ training sequences for PTB, and one canary per $\sim{\{3.5M,350K,70K,35K\}}$ training sequences for Wikitext.
V-B Training and Attack Implementation
FL Training Setup: We fine-tune Gemma-2b [team2024gemma], Llama-2 [touvron2023llama], GPT-2 base [radford2019language] and BERT [devlin-etal-2019-bert] as representative language models to evaluate our proposed attacks. Later we also observe the changes in attack performances for the medium and large versions of GPT-2. Due to computation resource limitations, we could not train the Llama-2 model in a large-scale FL setting. Instead, we conducted our attacks on Llama-2 in a compact experimental setup. Relevant experiments with Llama-2 are included in Appendix E. Additional information about these models is available in the Appendix B. To make fine-tuning large LMs feasible in FL, we adopt Parameter-Efficient Fine-Tuning (PEFT) [hu2022lora] for all models exceeding 1B parameters. Specifically, for Gemma-2b and Llama-2, we use QLoRA configurations consistent with recent large-scale FL and PEFT benchmarks. We apply LoRA to the attention projection matrices and MLP components. Unless otherwise specified, we use LoRA rank $r=16$, scaling factor $\alpha=32$, and dropout $p=0.05$, with trainable parameters restricted to LoRA adapters while keeping the base weights frozen. Following Dettmers et al. [dettmers2023qlora], we quantize the base model to 4-bit NF4.
@@ V-B Training and Attack Implementation
For the primary FL task, we have 100 participants and randomly select 20 for each round. We vary the number of attackers and victims between 1 and 10 and distribute the sensitive sequences among the victims accordingly. We train the model for 700 iterations, each with five local epochs, using an initial learning rate of $10^{-4}$ with a linearly decayed learning rate scheduler.
@@
❏We use a dataset of English plaintext jokes [pungas] as $\mathcal{D}_{reg}$ as discussed in Sections III and IV. This dataset exclusively contains $1622$ regular English sentences without privacy-sensitive sentences. Besides, we generate $100$ dummy privacy-sensitive sentences to use as $\mathcal{D}_{sen}.$ As mentioned earlier, it does not overlap with the original private texts. The attacker uses their prior knowledge of the sensitive data pattern to generate these dummy sentences and capture the model’s behavior towards out-of-distribution data. The number of epochs for finetuning $G_{r}$ and $G_{s}$ was empirically set to 20 with early stopping enabled.
@@
❏We use the Adam [kingma2014adam] optimizer in the selective weights optimization algorithm. We empirically tune the number of steps between 200 and 1000, the clipping threshold between 1 and 4, and set the initial learning rate at $1\times 10^{-3}$.
❏Since model weights do not fall into conventional categories of data, we experiment with several learning algorithms for the transformation task to identify the most suitable one. These algorithms include normalizing flows (e.g., planar flows [Lai-planner-flow] and RealNVP[dinh2016realNVP]), dense autoencoder (DAE), variational autoencoder (VAE)[kingma2013VAE], convolutional autoencoder (CAE)[zhang2018CAE], and transformer[vaswani_etal]. We find that DAE, CAE, and transformer showed comparable performances (in terms of validation loss) while normalizing flows and VAE did not yield promising outcomes. However, we eventually chose to employ DAE for this task as the other two, particularly the transformer, turned out to be computationally much more expensive taking roughly twice the time of DAE to complete the training.
❏To ensure better stability during the transformation, we use Z-score normalization [al2018effect] to normalize the weight vector before passing it to the SWO or WTL algorithm. After the transformation, we denormalize the output weight to its original scale. For the sake of simplicity, we skip the normalization step in Algorithms 1 and 2.
Evaluation Criteria: For the autoregressive text generation, we consider every sensitive sequence $x=p||c$ in the model’s training dataset, where x is split into a prefix $p$ (i.e., the non-sensitive component) and a suffix $c$ (i.e., the sensitive component). The suffix is the attacker’s point of interest. For each sequence, we report if the model exactly reproduces $c$ when prompted with $p$, following the data reconstruction procedure described in Section IV-F1. In masked language modeling, for every sensitive sequence x=pcsx=p||c||s in the training dataset, we aim to reconstruct the masked-out sensitive component $c$, given a prefix $p$ and suffix $s$. We also evaluate the sentence-level membership inference [mireshghallah2022quantifying] where the attacker attempts to infer whether a sensitive sequence $x$ is part of the victim’s local training data or not.
V-C Baselines and Metrics
@@
Intermediate Rounds (IR): The attacker has access to the ground truth information of intermediate victim rounds from server cooperation as mentioned in Section IV-D. IR directly prompts these victim models without applying any attack strategy.
@@
Metrics. We report the number of successful reconstructions (NSR) of the artificial canaries (for Wikitext and PTB) and the inherent sensitive texts (for Enron). We also evaluate the level of leakage using exposure (EXP), a metric introduced in [carlini2019secret], where higher exposure indicates greater privacy leakage. For membership inference, we report the recall and precision following [mireshghallah2022quantifying]. Moreover, we use validation perplexity (PPX) as a metric for the performance of the model, where lower perplexity refers to better model utility.
Figure 5: Number of successful reconstructions out of 200 unique canaries for different attack strategies along with the baselines for (a) Wikitext with Gemma, (b) Wikitext with GPT-2, (c) Wikitext with BERT. Then, out of 125 in-house sensitive sequences for (d) Enron with Gemma, (e) Enron with GPT-2, and (f) Enron with BERT. Finally, membership inference (g) recall and (h) precision scores for different attack strategies along with three baselines on the Wikitext Dataset using BERT.
VI Evaluation
We aim to answer the following research questions through a comprehensive empirical evaluation.
@@ VI Evaluation
RQ 4: How does the inclusion of the server as a passive adversary impact the attack performance?
VI-A Results
@@
TABLE II: Recall (%) and precision (%) of victim round identification step from 700 rounds.
VI-A1 Victim Round Identification Performance
Table II demonstrates the results of our victim round identification mechanism. We ablate the total number of victims from 1 to 10, and consider both static and dynamic attacks. With increased victim count and in dynamic attack mode, victim model updates come more frequently during training. It bolsters the global model’s memorization of the victim’s data, which in turn, helps in identifying the victim rounds. However, those extra victim updates during training impact non-victim training rounds as well, which generates more false-positive outputs. Hence, with the increased number of victims, the precision scores fall slightly (See confusion matrices in Figure 11). Nevertheless, our VRI method yields high precision and recall even in the most challenging setting of only a single victim, suggesting that it can identify the victim rounds with significant success, which answers RQ 1.
VI-A2 Data Reconstruction Performance
Figure 5(b) and 5(a) demonstrate the results across different attack strategies and the four baselines on the Wikitext dataset, in terms of the number of successful reconstructions when autoregressive text generation is performed with Gemma and GPT-2 respectively. Overall, The Gemma model achieves a superior NSR compared to GPT-2, mainly due to its larger size and complexity. First, comparing the scores of the first two baselines answers RQ 2. A higher score of IR compared to LR indicates that model snapshots from intermediate victim rounds cause greater leakage than only the final model (hypothesis 2), which justifies the importance of our victim round identification step. Next, for both Gemma and GPT-2, our static attack outperforms IR by 15-18% with both SWO and WTL, answering RQ 3. This shows that it is indeed possible to significantly increase the LM’s memorization of clients’ private data by maliciously tampering with its weights. Additionally, WTL performs better than SWO in all cases, proving it to be a better MDM technique. Apart from that, introducing active threats during training further increases privacy leakage, as our dynamic mode of the attack has a much better NSR (137 for Gemma) than the static mode (125). Furthermore, incorporating victims’ local updates in place of their global counterparts with cooperation from the server boosts the attack performance considerably (by 6 to 8%) in both $static^{+}$ and $dynamic^{+}$ modes of the attack. In fact, $dynamic^{+}$ is our best-performing attack variant with $71\%$ (NSR=142) successful reconstructions.
@@ VI-A2 Data Reconstruction Performance
Figure 5(c) illustrates attack performance in masked language modeling, i.e., with BERT. Overall, NSR scores for BERT are much lower than GPT-2. The improvement in attack performance for WTL over SWO is also not as prominent in this case as it was for GPT-2. Jagannatha et al. [jagannatha2021membership] show that masked LMs have lower privacy leakage than autoregressive LMs. This could be attributed to the difference in the architecture of these two types of LMs and how they understand the context. BERT understands the context from both the left and the right sides of a masked-out word, whereas Gemma/ GPT-2 uses only the left context to predict the next word. BERT’s bidirectional attention might help it distribute its focus, thereby reducing the chances of overfitting to specific examples. However, our best attack strategy without any server cooperation (dynamic) yields around 95 out of 200 successful reconstructions, while with $dynamic^{+}$the NSR reaches $55\%$. From the results, it is quite evident that our proposed attack methods are highly effective against masked language modeling in addition to autoregressive text generation. Due to space constraints, we put the results on the PTB dataset in Appendix F.
@@
Finally, the performance of our attacks on real-world data, i.e., the Enron email dataset, is depicted in Figures 5(e) and 5(f). It is worth mentioning that the privacy-sensitive sequences in this dataset are less structured than the artificial canaries, making the attacks more challenging and having some interesting impact on the results. Firstly, both the static and $static^{+}$ attacks slightly underperform in this scenario compared to artificial canary extraction. As shown in Figures 5(e) and 5(f), the NSR scores of the static attack for both Gemma/GPT-2 and BERT are quite close to IR’s score. Nevertheless, active interference by the adversary changes the picture notably, as we can see the dynamic attack on Gemma increases the NSR by around 10% from the static variant, while the $dynamic^{+}$ attack manages to extract 64% (80 out of 125) of the private sequences. The results show that our attacks can successfully compromise the privacy of real-world data with high accuracy. Apart from that, similar to the Wikitext dataset with canaries, masked language modeling with BERT turns out to be more resilient than Gemma/GPT-2 against leakage for the Enron dataset with real-world private data (Figure 5(f)). Results with respect to the EXP metric are moved to the Appendix (D).
VI-A3 Membership Inference Performance
@@
We adopt the implementation of [mireshghallah2022quantifying] for membership inference based on the likelihood ratio hypothesis, where part of our test set consists of all the canaries (for PTB and Wikitext) or private texts (for Enron). We insert 200 canary-like false-positive (outside training data) samples into the test set, which are semantically close to the canaries but do not match any of their secret portions. The evaluation is done over 400 samples (200 canaries + 200 false positives) for both Wikitext and PTB. For the Enron dataset, we add 125 false positives into the test set along with the 125 already identified private sequences, making 250 samples in total. Looking at the results for Wikitext in Figure 5(g), we can see that the static mode significantly improves the recall score (by 15-26%) compared to the lower baselines (LR, IR). This demonstrates the effectiveness of our MDM algorithms in identifying the canaries and confidently inferring their membership. We achieve the highest recall of 61% with the $dynamic^{+}$ version, which shows results similar to the upper baselines (FP and MU). Our MDM approach directs the language model to prioritize sensitive training samples, leading to assigning a higher likelihood to those samples. This reduces the number of missed positive samples, increasing the true positive count and decreasing the false negative count. Consequently, the recall rate experiences a significant improvement due to these dual effects. However, there was only a slight improvement in the ratio of true positives to false positives. As a result, the attack performance did not fluctuate much across different attacks and the baselines in terms of precision as shown in 5(h). For a better understanding of the phenomenon, please refer to the confusion matrices in Figure 11 and Section G-C of the Appendix. We also put the membership inference results for PTB and Enron into the Appendix (Section G-B) for space constraints.
We conducted extensive ablation studies on our attack performance based on six different criteria: data repetitions, prompt perturbation, number of sensitive layers, number of victims and attackers, frequency of victim rounds, and model size. Please refer to Appendix I for details.
VI-B Attack Cost Analysis